OpenStudy (anonymous):

int_{0}^{3} x/(sqrt{x^2+16}

OpenStudy (anonymous):

let\[x =\tan \theta\]

OpenStudy (anonymous):

:) thx

OpenStudy (anonymous):

On closer inspection, you don't need trig sub, u sub would suffice. I never noticed until after I worked it out myself. But hey, good practice for trig sub. When you get it on the test you going to knock it out of the park.

OpenStudy (amistre64):

shouldnt x = 4tan(t)? that way; when you square it you get: sqrt(16 tan^2(t) + 16) = sqrt(16(tan^2+1))) = 4sec^2

OpenStudy (amistre64):

err.... 4 sec :)

OpenStudy (anonymous):

Yes. I think he picked that up.